### v2

# Reliability entails input-selective contraction and regulation in excitable networks

Abstract
The animal nervous system offers a model of computation combining digital reliability and analog efficiency. Understanding how this sweet spot can be realized is a core question of neuromorphic engineering. To this aim, this paper explores the connection between reliability, contraction, and regulation in excitable systems. Using the FitzHugh-Nagumo model of excitable behavior as a proof-of-concept, it is shown that neuronal reliability can be formalized as an average trajectory contraction property induced by the input. In excitable networks, reliability is shown to enable regulation of the network to a robustly stable steady state. It is thus posited that regulation provides a notion of dynamical analog computation, and that stability makes such a computation model robust.
I Introduction
Neurons are the archetype of excitable dynamics; they respond to some input stimuli with sudden, short-lived spikes that carry asynchronous, digital information atop the neuron’s continuous trajectories. The animal nervous system is the archetype of excitable control system; while boasting the efficiency of analog, sparse computation, it gives rise to reliable control and information processing without relying on clocks or global coordination. Understanding how reliable computation emerges from the analog dynamics of excitable systems is a central question of neuromorphic control [24, 27, 26]; its answer is indeed key to devise a principled design theory for neuromorphic systems, which is currently missing.
This paper approaches such a question by starting from the neuroscience notion of reliability, a basic property of biological neurons by which they respond to some specific inputs with a spike pattern that, except for an initial transitory, does not depend on the specific state of the neuron before the input is applied [20, 15, 8, 13]. In the simplifying, yet emblematic context of FitzHugh-Nagumo (FHN) models, this paper develops a new mathematical theory of reliability based on contraction analysis [19, 6]. In particular, it is shown that the state space of an excitable system is divided into three regions: two contraction regions separated by a third “expansion region” where contraction does not hold. The system’s rest point lies in one of the two contraction regions; each spike brings the system’s state to the other and then back to the first. In-between, the state passes through the expansion region. As a consequence, all inputs keeping the trajectories within the contraction regions for enough time will induce, on average, contraction of the trajectories. Contraction, in turn, is what leads to a reliable behavior as it implies that the initial conditions are forgotten as time flows.
Reliability is a well-studied property in the computational neuroscience literature, and it is the subject of several existing theoretical and numerical studies [17, 25, 4, 29, 12, 28]. However, these mostly focus on noisy inputs, which are not a relevant class to characterize deterministic computation and control, and their theoretical machinery is based on Lyapunov exponents, which are theoretically unsatisfactory for several reasons: negative Lyapunov exponents do not actually guarantee stability of trajectories [18]; they only give information on the system around the trajectory on which they are computed; they can only be numerically estimated after simulating a trajectory for a long time; (as a consequence) they provide an information that cannot be collocated in time.
Unlike existing approaches, the theory developed in this paper provides mathematical guarantees and works for arbitrary inputs. In addition, and more importantly, characterizing reliability in terms of contraction creates a link between reliability and computation. In particular, it is well-known in the output regulation literature [10, 23, 22, 2, 11] that, within a control system, closed-loop contraction implies robust regulation: all trajectories converge to a unique, robustly stable steady state where the regulation objectives are met. In the same way, in excitable networks reliability yields regulation as it implies contraction. In turn, regulation provides a notion of dynamical computation: if a network regulates when subject to some input, the steady-state trajectory to which all trajectories converge can be interpreted as the result of the network’s computation for that specific input. Turing’s classical model of computation can be obtained as a special case when both the input and the corresponding response are constant in time; otherwise, computation assumes a dynamical character. As regulation entails robust stability of the steady state, this notion of computation is robust (robustness can also be enhanced further by focusing on the specific events present in the trajectories – spikes, bursts, etc. – instead of considering their full time behavior [26]).
As a second objective, this paper develops such a connection between excitability, regulation, and computation by extending the reliability analysis to a benchmark excitable network recently proposed in [7], which exhibits homeostasis when driven by a specific class of inputs. In this context, owing to the regulation analogy, computation is linked to the theory of internal models [3].
II Excitability, Contraction, and Reliability
II-A Contraction Notions
The notion of contraction used in this paper is adapted from [19, 6]. A system $\dot{x}=f(x,t)$, with state $x(t)\in\mathbb{R}^{n}$, is said to be contractive on a convex set $C\subset\mathbb{R}^{n}$ with respect to a metric $d:\operatorname{dom}d\to[0,\infty)$, where $\operatorname{dom}d\supset C\times C$, if there exists $\lambda>0$ such that, for every $t_{0}\in\mathbb{R}$, every two solutions $x_{a}$ and $x_{b}$ of the system satisfy $d(x_{a}(t),x_{b}(t))\leq e^{-\lambda(t-t_{0})}d(x_{a}(t_{0}),x_{b}(t_{0}))$ for every $t\geq t_{0}$ for which $x_{a}(\tau)\in C$ and $x_{b}(\tau)\in C$ for all $\tau\in[t_{0},t]$. The set $C$ is called a contraction region and the constant $\lambda$ the contraction rate. In addition, given two solutions $x_{a}$ and $x_{b}$, two time instants $t_{0},t_{1}\in\mathbb{R}$, and a scalar $\alpha\in[0,1)$, the solutions $x_{a}$ and $x_{b}$ are said to $\alpha$-contract from $t_{0}$ to $t_{1}$ with respect to $d$ if $d(x_{a}(t_{1}),x_{b}(t_{1}))\leq\alpha d(x_{a}(t_{0}),x_{b}(t_{0}))$. They are said to contract from $t_{0}$ to $t_{1}$ with respect to $d$ if $d(x_{a}(t_{1}),x_{b}(t_{1}))<d(x_{a}(t_{0}),x_{b}(t_{0}))$. Contraction formalizes reliability: if the dependency on time in the system is due to a fixed external input, and different initial conditions model the application of such an input to the same system at different times, then contraction implies asymptotic agreement across trials.
II-B Input-Selective Contraction of FitzHugh-Nagumo Models
Consider the following FHN system [9]
$\dot{v}$ $=v-\tfrac{1}{3}v^{3}-w+u,$ $\dot{w}$ $=\varepsilon(v-bw+a)$ (1)
with state $x=(v,w)\in\mathbb{R}^{2}$, and where
$1-2b/3<a<1,\quad b\in(0,1),\quad\varepsilon<1/b.$ (2)
In the neuronal context, $v$ represents the membrane potential of the neuron, $w$ represents a slow dynamics recovering stability after the spike, and $u$ represents the effect of the input stimulus on the membrane potential. For any $\mu>0$, define the sets $C_{\downarrow}(\mu)\coloneqq\{x\in\mathbb{R}^{2}\mid v\leq-\sqrt{1+\mu}\}$ and $C^{\uparrow}(\mu)\coloneqq\{x\in\mathbb{R}^{2}\mid v\geq\sqrt{1+\mu}\}$. Both are convex and closed. Moreover, with $\varepsilon$ the same as in (1), define the following metric on $\mathbb{R}^{2}$
$d(x_{a},x_{b})\coloneqq\sqrt{\tfrac{1}{2}(v_{a}-v_{b})^{2}+\tfrac{1}{2\varepsilon}(w_{a}-w_{b})^{2}}.$ (3)
Then, the following result shows that, for every $\mu>0$, both $C_{\downarrow}(\mu)$ and $C^{\uparrow}(\mu)$ are contraction regions.
Proposition 1.
For every $\mu>0$, system (1) is contractive on both $C_{\downarrow}(\mu)$ and $C^{\uparrow}(\mu)$ with respect to the metric (3) and with contraction rate $\lambda(\mu)=\min\{\mu,b\varepsilon\}$.
Proof.
The Jacobian of (1) is
$J(x)=\begin{bmatrix}1-v^{2}&-1\\ \varepsilon&-b\varepsilon\end{bmatrix}$ (4)
Moreover, $d$ satisfies $d(x_{a},x_{b})^{2}=(x_{a}-x_{b})^{\top}P(x_{a}-x_{b})$ with $P\coloneqq\frac{1}{2}\operatorname{diag}(1,\ \varepsilon^{-1})$. Then, for every $\mu>0$ and every $x\in C^{\uparrow}(\mu)\cup C_{\downarrow}(\mu)$, $J(x)^{\top}P+P^{\top}J(x)=\operatorname{diag}(1-v^{2},-b)\leq-\operatorname{diag}(\mu,b)$. The claim then follows from this latter inequality and from the convexity of $C^{\uparrow}(\mu)$ and $C_{\downarrow}(\mu)$ by means of well-known arguments based on Hadamard’s Lemma; see, e.g., [6, Lem 3.1, Thm. 3.9].
Proposition 1 implies that, for every $\mu>0$, every two solutions of (1) converge to one another in the metric (3) at exponential rate $\lambda(\mu)$ on any interval of time in which they lie in the same region $C_{\downarrow}(\mu)$ or $C^{\uparrow}(\mu)$. Taking the union for all $\mu>0$ leads to the definition of the following sets
$C_{\downarrow}$ $\coloneqq\{x\in\mathbb{R}^{2}\mid v<-1\},$ $C^{\uparrow}$ $\coloneqq\{x\in\mathbb{R}^{2}\mid v>1\},$
which we call the lower and upper contraction region, respectively. Clearly, if $x\in C^{\uparrow}$ (resp. $C_{\downarrow}$), then it is in $C^{\uparrow}(\mu)$ (resp. $C_{\downarrow}(\mu)$) for all small enough $\mu>0$. Hence, as long as two trajectories move together inside $C^{\uparrow}$ or $C_{\downarrow}$, their distance decreases with time. However, it is important to notice that such a decrease may not happen at an exponential rate, as the contraction rate vanishes if one of the two solutions approaches the boundary of $C^{\uparrow}\cup C^{\downarrow}$. The complement of the union of the two contraction regions, namely $I\coloneqq\mathbb{R}^{2}\setminus(C_{\downarrow}\cup C^{\uparrow})=\{x\in\mathbb{R}^{2}\mid v^{2}\leq 1\}$, defines a closed set where the FHN model (1) is not contractive. This region is therefore called the expansion region, and it separates the two contraction regions.
For $u=0$, System (1) has an equilibrium $x^{\star}=(v^{\star},w^{\star})$ given by the real solution of $v^{3}+\frac{3}{b}(1-b)v+3\frac{a}{b}=0$ and vbw+a=0v-bw+a=0. For the range of parameters (2), one has $v^{\star}<-1$. Indeed, assume ad absurdum that $v^{\star}\geq-1$. Then, the equilibrium conditions and (2) imply $v^{\star}\in[-1,0]$ and, hence, $1\leq(v^{\star})^{3}=-\tfrac{3}{b}(1-b)v^{\star}-3\frac{a}{b}\leq\tfrac{3}{b}(1-b)-3\frac{a}{b}<-1$, a contradiction. With similar arguments, it can also be shown that a spike-inducing stimulus will bring $v$ to values larger than $1$. Thus, the stable equilibrium $x^{\star}$ of the FHN model lies in the lower contraction region $C_{\downarrow}$ and every spike brings the state to the upper contraction region $C^{\uparrow}$. A spike-inducing input has therefore the effect of making the system switch between the two contraction regions, from the lower to the upper, and back. In-between, the system crosses the expansion region $I$ where the trajectories may diverge to one another. In the time-scale of more spikes, the overall contraction property of the system depends on the relative time spent in the two contraction regions. In particular, let $x_{a}$ and $x_{b}$ be any two solutions of (1) corresponding to some common input $u$. Pick two time instants $t_{0},t_{1}\in\mathbb{R}$ such that $t_{1}\geq t_{0}$, and let $T_{\rm c}\subset[t_{0},t_{1}]$ denote the set of time instants $t$ where $x_{a}(t)$ and $x_{b}(t)$ are both in $C^{\uparrow}$ or both in $C_{\downarrow}$. Finally, let $T_{\rm i}\coloneq[t_{0},t_{1}]\setminus T_{\rm c}$. Since
$\int_{0}^{1}J(sx_{a}+(1-s)x_{b}){\rm d}s=\begin{bmatrix}1-\frac{1}{3}(v_{a}^{2}+v_{b}^{2}+v_{a}v_{b})&-1\\ \varepsilon&-b\varepsilon\end{bmatrix}$,
then Hadamard’s Lemma implies
$\frac{{\rm d}}{{\rm d}t}d(x_{a}(t),x_{b}(t))=-\nu(x_{a}(t),x_{b}(t))+3$ (5)
for all $t\in[t_{0},t_{1}]$, with $\nu(x_{a},x_{b})\coloneq 3+\frac{1}{3}(v_{a}^{2}+v_{b}^{2}+v_{a}v_{b}-3)(v_{a}-v_{b})^{2}+b(w_{a}-w_{b})^{2}$. Hence, the solutions $x_{a}$ and $x_{b}$ contract from $t_{0}$ to $t_{1}$ with respect to $d$ if and only if
$\int_{T_{\rm c}}\nu(x_{a}(t),x_{b}(t)){\rm d}t>3(t_{1}-t_{0})-\int_{T_{\rm i}}\nu(x_{a}(t),x_{b}(t)){\rm d}t.$
From this relation, we can derive a simpler sufficient criterion for $\alpha$-contraction of a group of trajectories that only relies on the average time spent by the trajectories in the same contraction region. Consider an input $u$, a set $X_{0}\subset\mathbb{R}^{2}$ of initial conditions for (1), and two time instants $t_{0},t_{1}\in\mathbb{R}$ such that $t_{1}\geq t_{0}$. Let $\mathcal{S}(X_{0},u,t_{0})$ denote the set of solutions of (1) corresponding to the input $u$ and starting in $X_{0}$ at time $t_{0}$. Pick $\mu>0$ arbitrarily and let $\Delta_{\rm c}(\mu)$ denote the total amount of time spent by all the considered trajectories, together, in $C^{\uparrow}(\mu)$ or in $C_{\downarrow}(\mu)$. Then, the following holds.
Proposition 2.
With $\lambda(\mu)=\min\{\mu,b\varepsilon\}$, let $\alpha\coloneq e^{-\lambda(\mu)\Delta_{\rm c}(\mu)-\Delta_{\rm c}(\mu)+t_{1}-t_{0}}$. If $\Delta_{\rm c}(\mu)>(t_{1}-t_{0})/(1+\lambda(\mu))$, then $\alpha\in[0,1)$ and any two solutions in $\mathcal{S}(X_{0},u,t_{0})$ $\alpha$-contract from $t_{0}$ to $t_{1}$ with respect to the metric (3).
Proof.
Let $T_{\rm c}(\mu)$ denote the set of time instant in which all solutions in $\mathcal{S}(X_{0},u,t_{0})$ are either all in $C^{\uparrow}(\mu)$ or all in $C_{\downarrow}(\mu)$. Pick any two solutions $x_{a},x_{b}\in\mathcal{S}(X_{0},u,t_{0})$. In view of Proposition 1, for every $\tau,s\in\mathbb{R}$ such that $(\tau,s)\subset T_{\rm c}(\mu)$, $d(x_{a}(s),x_{b}(s))\leq e^{-\lambda(\mu)(s-\tau)}d(x_{a}(\tau),x_{b}(\tau))$. Moreover, in view of (5), and since $v_{a}^{2}+v_{b}^{2}+v_{a}v_{b}\geq 0$ for all $v_{a},v_{b}\in\mathbb{R}$, for every $t\in[t_{0},t_{1}]\setminus T_{\rm c}(\mu)$ we obtain $\frac{\rm d}{{\rm d}t}d(x_{a}(t),x_{b}(t))^{2}\leq 2d(x_{a}(t),x_{b}(t))^{2}$. Grönwall’s Lemma, thus implies $d(x_{a}(t_{1}),x_{b}(t_{1}))\leq e^{-\lambda(\mu)\Delta_{\rm c}(\mu)+t_{1}-t_{0}-\Delta_{\rm c}(\mu)}d(x_{a}(t_{0}),x_{b}(t_{0}))$, from which the results follows.
A
B
C
D
Figure 1: Simulations of (1) with $a=0.7$, $b=0.8$, and $\varepsilon=1/12.5$ from $9$ different initial conditions. Light-blue patches highlight the time intervals where all solutions lie within the same contraction region; (top) the solid, dark-blue lines depict the time evolution of the $v$ variable for each of the nine solutions; dashed red lines depict the boundaries of the contraction regions; (center) pointwise maximum distance between the solutions using the metric (3); (bottom) applied input $u$.
II-C Contraction and Reliability
This section illustrates on four benchmark scenarios how the contraction analysis developed in the previous section can be used to explain the input-selective reliability behavior observed in biological neurons. The first scenario considers a constant input $u(t)=0.7$. When a large-enough constant input is given to a neuron, it leads to a sustained periodic train of spikes. Solutions originating from different initial conditions do not contract, but keep a constant phase shift. This behavior is well-documented experimentally [20, 15] and is replicated by the FHN model (1): Fig. 1-A shows the evolution of $v(t)$ subject to $u(t)=0.7$ for several different initial conditions around the stable equilibrium. From the viewpoint of Section II-B, this behavior is explained by the fact that the effect of contraction and instability balance out in the time scale of multiple spikes, leading to no appreciable trend in the time evolution of the distance between solutions. In the second scenario, a small harmonics with a frequency similar to the natural frequency of oscillation of the FHN model in the previous case is added to the previous constant stimulus; the resulting input is $u(t)=0.7+0.2\sin(\pi t/23)$. The obtained behavior is depicted in Fig. 1-B. Consistently to the experimental evidence [15], we now observe a reliable behavior due to a resonance effect that leads to slightly sharper spikes spending more time in the contraction regions. The third scenario uses as input a slow square pattern of period $60$, amplitude $0.6$, and duty-cycle $1/3$. The observed behavior, depicted in Figure 1-C, is reliable. Reliability follows because the forcing period is significantly larger than the natural one; hence, the state spends most time close to its equilibrium in the lower contraction region. Average contraction is then a consequence of canonical dwell-time results in switched systems [14]. Finally, the last scenario simulates the response to a Gaussian noisy input whose bias and variance change in time: in the interval $[0,100)$, the bias equals $0.5$ and the variance $0.25$; in the interval $[100,200)$, the bias is $0.3$ and the variance $0.01$; for the remaining time, the bias is $0.1$ and the variance $0.0025$. The simulation results shown in Fig. 1-D are consistent with the experimental evidence [20, 15, 12] and can be explained in terms of average contraction: high noise contains non-negligible power around the resonance frequency [15]; as in the second scenario, resonance fosters reliability (the distance between trajectories decreases in the interval $[0,100)$); medium noise induces an unreliable behavior (the distance between trajectories increases in the interval $[100,200)$) because the input is powerful enough to bring the solutions outside the lower contraction region but not powerful enough to make them consistently visit the upper contraction region; finally, small noise leads again to a reliable behavior since the input is not powerful enough to bring the trajectories out of the lower contraction region. This is an example of subthreshold reliability.
III Reliability Enables Regulation
The motif depicted in Fig. 2-A, to which we will refer to as the EI network, provides a simple, yet emblematic example of how reliability enables regulation in excitable networks. The EI network is a basic excitatory-inhibitory motif constructed in [7] in solving the following event disturbance rejection problem. Consider an excitable controlled system ($\rm E$ in Fig. 2-A) described by the following FHN model
${\rm E}\!:\!\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\left\{\normalcolor\begin{aligned} \dot{v}_{\text{\tiny E}}&=v_{\text{\tiny E}}-\tfrac{1}{3}v_{\text{\tiny E}}^{3}-w_{\text{\tiny E}}+k_{\text{\tiny E}}(u-\eta+V_{\rm rest}-y)\\ \dot{w}_{\text{\tiny E}}&=\varepsilon_{\text{\tiny E}}(v_{\text{\tiny E}}-b_{\text{\tiny E}}w_{\text{\tiny E}}+a_{\text{\tiny E}})\\ y&=v_{\text{\tiny E}}\end{aligned}\right$. (6)
in which $V_{\rm rest}$ denotes the rest point of the variable $v_{\text{\tiny E}}$ in the absence of coupling (i.e., with $k_{\text{\tiny E}}=0$), and the term $k_{\text{\tiny E}}(u-\eta+V_{\rm rest}-y)$ represents a diffusive coupling with an excitatory (entering with a plus) input $u$ representing a disturbance, and an inhibitory (entering with a minus) input $\eta$ representing a control input. The coupling gain $k_{\text{\tiny E}}$ is a control parameter. The control goal is to design a controller, made of excitable systems, to reject the disturbance $u$ through the control input $\eta$. This is an output regulation problem that is solvable under the assumption that the disturbance $u$ is generated by a known system, called the exosystem. In [7], the exosystem is taken as a simple half-center oscillator [21] depicted in Fig. 2-B and described by the following equations
${\rm U}\!:\!\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\left\{\normalcolor\begin{aligned} \dot{v}_{\text{\tiny U}1}&=v_{\text{\tiny U}1}-\tfrac{1}{3}v_{\text{\tiny U}1}^{3}-w_{\text{\tiny U}1}+k_{\text{\tiny U}}(-v_{\text{\tiny U}2}-v_{\text{\tiny U}1})\\ \dot{w}_{\text{\tiny U}1}&=\varepsilon_{\text{\tiny U}}(v_{\text{\tiny U}1}-b_{\text{\tiny U}}w_{\text{\tiny U}1}+a_{\text{\tiny U}})\\ \dot{v}_{\text{\tiny U}2}&=v_{\text{\tiny U}2}-\tfrac{1}{3}v_{\text{\tiny U}2}^{3}-w_{\text{\tiny U}2}+k_{\text{\tiny U}}(-v_{\text{\tiny U}1}-v_{\text{\tiny U}2})\\ \dot{w}_{\text{\tiny U}2}&=\varepsilon_{\text{\tiny U}}(v_{\text{\tiny U}2}-b_{\text{\tiny U}}w_{\text{\tiny U}1}+a_{\text{\tiny U}})\\ u&=v_{\text{\tiny U}1}.\end{aligned}\right$. (7)
In particular, $\rm U$ is given by the reciprocal inhibitory diffusive interconnection of two FHN models of the form (1). Under the assumption that $u$ is generated by the exosystem $\rm U$, the internal model principle [3] suggests that a controller solving the problem must embed an internal model of the exosystem. The solution proposed in [7] is a minimal network with such a property. It consists of a controller realized by the block $\rm I$ in Fig. 3-A, described by the following equations
${\rm I}\!:\!\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\left\{\normalcolor\begin{aligned} \dot{v}_{\text{\tiny I}1}&=v_{\text{\tiny I}1}-\tfrac{1}{3}v_{\text{\tiny I}1}^{3}-w_{\text{\tiny I}1}+k_{\text{\tiny U}}(y-V_{\rm rest}-v_{\text{\tiny I}2}-v_{\text{\tiny I}1})\\ \dot{w}_{\text{\tiny I}1}&=\varepsilon_{\text{\tiny U}}(v_{\text{\tiny I}1}-b_{\text{\tiny U}}w_{\text{\tiny I}1}+a_{\text{\tiny U}})\\ \dot{v}_{\text{\tiny I}2}&=v_{\text{\tiny I}2}-\tfrac{1}{3}v_{\text{\tiny I}2}^{3}-w_{\text{\tiny I}2}+k_{\text{\tiny U}}(-v_{\text{\tiny I}1}-v_{\text{\tiny I}2})\\ \dot{w}_{\text{\tiny I}2}&=\varepsilon_{\text{\tiny U}}(v_{\text{\tiny I}2}-b_{\text{\tiny U}}w_{\text{\tiny I}1}+a_{\text{\tiny U}}).\end{aligned}\right$. (8)
When $y=V_{\rm rest}$, system (8) equals the exosystem (7). Hence, $\rm I$ embeds an internal model of $\rm U$. As a direct consequence of such an internal model property, for every disturbance $u$ generated by (7), there exists an initialization for the EI network (6), (8) – namely, the state of $\rm E$ at rest and that of $\rm I$ matching the initialization of $\rm U$ producing $u$ – leading to a steady-state trajectory along which $y(t)=V_{\rm rest}$ for all $t\in\mathbb{R}$, i.e., along which the disturbance $u$ is completely rejected. The considered problem is then solved if such a target steady state is attractive. This is possible if the EI network (6), (8) is contractive. However, as noted in [7], forcing contraction by a specific choice of the control parameter $k_{\text{\tiny E}}$ is difficult to imagine. Instead, we posit that, as for reliability in single neurons, contraction should rather be conceived here as a property induced by the input, which in this case is the disturbance $u$ itself: the solution of the regulation problem at hand turns out to be a question of network reliability.
A
B
Figure 2: (A) The EI network; (B) the exosystem.
Studying the reliability of the EI network requires a generalization of Proposition 1. Such a generalization is given by the following result, which shows that, similarly to the single-neuron case, also the EI network admits a decomposition of the state space (now $6$-dimensional) in contraction regions separated by expansion regions. Let $\sigma$ be the maximum eigenvalue of the matrix
$K\coloneqq\begin{bmatrix}-k_{\text{\tiny E}}&\frac{1}{2}(k_{\text{\tiny U}}-k_{\text{\tiny E}})&0\\ \frac{1}{2}(k_{\text{\tiny U}}-k_{\text{\tiny E}})&-k_{\text{\tiny U}}&-k_{\text{\tiny U}}\\ 0&-k_{\text{\tiny U}}&-k_{\text{\tiny U}}\end{bmatrix}$. (9)
Notice that $\sigma\geq 0$ since $\rho=\begin{bmatrix}0&1&-1\end{bmatrix}$ satisfies $\rho K\rho^{\top}=0$. Moreover, with $P\coloneqq\operatorname{diag}(1,\varepsilon_{\text{\tiny E}}^{-1},1,\varepsilon_{\text{\tiny U}}^{-1},1,\varepsilon_{\text{\tiny U}}^{-1})$, define the following metric on $\mathbb{R}^{6}$
$d(x,x^{\prime})=\sqrt{(x-x^{\prime})^{\top}P(x-x^{\prime})},$ (10)
where $x$ is used as a short for $x=(v_{\text{\tiny E}},w_{\text{\tiny E}},v_{\text{\tiny I}1},w_{\text{\tiny I}1},v_{\text{\tiny I}2},w_{\text{\tiny I}2})$. Then, the following result extends Proposition 1 to the case of the EI network.
Proposition 3.
For every $\mu>\sigma$, the EI network (6), (8) is contractive on any convex subset of
$C(\mu)\coloneqq\big\{x\in\mathbb{R}^{6}\mid v_{\text{\tiny E}}^{2}\geq 1+\mu,\ v_{\text{\tiny I}1}^{2}\geq 1+\mu,\ v_{\text{\tiny I}2}^{2}\geq 1+\mu\big\}$
with respect to the metric (10).
Proof.
Let us compactly rewrite the EI network (6), (8) as $\dot{x}=f(x)$, with $x=(v_{\text{\tiny E}},w_{\text{\tiny E}},v_{\text{\tiny I}1},w_{\text{\tiny I}1},v_{\text{\tiny I}2},w_{\text{\tiny I}2})$. Change coordinates as $x\mapsto z=Tx$ where $T\in\mathbb{R}^{6\times 6}$ is the permutation matrix such that $z=(v_{\text{\tiny E}},v_{\text{\tiny I}1},v_{\text{\tiny I}2},w_{\text{\tiny E}},w_{\text{\tiny I}1},w_{\text{\tiny I}2})$. In the new variables, $\dot{z}=g(z)$ with $g(\cdot)\coloneqq Tf(T^{-1}\,\cdot\,)$. The Jacobian of $g$ reads
$J(z)=\begin{bmatrix}A(z)&-I\\ E&-EB\end{bmatrix}$
in which $E\coloneqq\operatorname{diag}(\varepsilon_{\text{\tiny E}},\varepsilon_{\text{\tiny E}},\varepsilon_{\text{\tiny E}})$, $B\coloneqq\operatorname{diag}(b_{\text{\tiny E}},b_{\text{\tiny E}},b_{\text{\tiny E}})$, and
$A(z)\coloneq\begin{bmatrix}1-v_{\text{\tiny E}}^{2}-k_{\text{\tiny E}}&-k_{\text{\tiny E}}&0\\ k_{\text{\tiny U}}&1-v_{\text{\tiny I}1}^{2}-k_{\text{\tiny U}}&-k_{\text{\tiny U}}\\ 0&-k_{\text{\tiny U}}&1-v_{\text{\tiny I}2}^{2}-k_{\text{\tiny U}}\end{bmatrix}$
Let $Q\coloneqq\frac{1}{2}\operatorname{diag}(I,\ E^{-1})$. Then, $J(z)^{\top}Q+QJ(z)=\operatorname{diag}(\tfrac{1}{2}(A(z)^{\top}+A(z)),\ -B)$. Since $\frac{1}{2}(A(z)^{\top}+A(z))=\operatorname{diag}(1-v_{\text{\tiny E}}^{2},1-v_{\text{\tiny I}1}^{2},1-v_{\text{\tiny I}2}^{2})+K$, where $K$ is defined in (9), then $v_{\text{\tiny E}}^{2}\geq 1+\mu$, $v_{\text{\tiny I}1}^{2}\geq 1+\mu$, and $v_{\text{\tiny I}2}^{2}\geq 1+\mu$ with $\mu>\sigma$ imply $\frac{1}{2}(A(z)^{\top}+A(z))\leq-(\mu-\sigma)I<0$. The claim then follows by means of the same arguments used in the proof of Proposition 1 and by noticing that $T^{\top}QT=P$.
A
B
C
Figure 3: Simulations of $10$ trials of the EI network (6), (8) with $a_{\text{\tiny E}}=0.7$, $b_{\text{\tiny E}}=0.8$, $\varepsilon_{\text{\tiny E}}=1/12.5$, $k_{\text{\tiny E}}=4$, $a\text{\tiny U}=0.6$, $b\text{\tiny U}=0.7$, $k\text{\tiny U}=1/2$, and $\varepsilon_{\text{\tiny U}}=1/30$. (A) $u$ is generated by the exosystem (7); (B), $u(t)=2\sin(\pi t/30)$; (C) $u$ is generated by (7) with $\varepsilon_{\text{\tiny U}}=1/5$ (while in (8), $\varepsilon_{\text{\tiny U}}$ remains $1/30$). In all figures: (top) the $10$ simulated trajectories of $y(t)$; (bottom) dark blue lines depict the $10$ simulated trajectories of the variable $\eta$, orange lines depict the disturbance $u$.
The reader will notice the similarity between $C(\mu)$ and $C_{\downarrow}(\mu)\cup C^{\uparrow}(\mu)$ of Section II-B. We also notice that, if $k_{\text{\tiny E}}=k_{\text{\tiny U}}$ (namely, when the network is balanced), then $\sigma=0$, and $\mu$ is only constrained to be strictly positive as in Proposition 1. In view of Proposition 3, we can conclude that, if the input $u$ makes the EI network’s state spend enough time in the contraction regions, then the EI network contracts and regulation occurs. All disturbances $u$ generated by the exosystem (7) belong to this regulation-inducing class. Indeed, for each $u$ produced by (7), along the corresponding target steady-state trajectory $y(t)=V_{\rm rest}<-1$ and the $\rm I$ system oscillates in phase with (7). As a consequence, we can conclude that the EI network structurally rejects disturbances produced by the exosystem (7) as these induce regulation.
Fig. 3-A illustrates the previous result; in response to disturbance $u$ generated by (7), in all trials $y(t)$ converges to $V_{\rm rest}$ (top) and $\eta(t)$ synchronizes with $u(t)$ (bottom). In addition, Fig. 3-B reports a simulation in which $u$ is a sinusoid of similar frequency and amplitude of the disturbance of Fig. 3-A. Although in this case perfect rejection is not possible, since the internal model differs from the generator of a sinusoid, contraction still takes place. Indeed, input-selective contraction is robust with respect to perturbations of the input. Moreover, we also observe that the events – defined as the peaks of the oscillating patterns – in $\eta$ and $u$ synchronize, and, consequently, the output trajectory $y$ does not spike. In the sense of events, the “perturbed” disturbance is therefore effectively filtered out as it induces no spike in $y$. This example illustrates the robustness of event regulation [26] contrasted to the fragility of exact trajectory regulation [2]. Lastly, Fig. 3-C shows a simulation with an exosystem that is modified by setting $\varepsilon_{\text{\tiny U}}=1/4$ while keeping $\varepsilon_{\text{\tiny U}}=1/30$ in (8). In this case, contraction does not hold any more; the trajectories of $\eta$ and $y$ do not synchronize across trials, and regulation does not take place.
Finally, we notice that, to compensate for $u$, $\eta$ must asymptotically match $u$. Hence, the EI network may also be interpreted as performing tracking of $u$ in the variable $\eta$. In this case, Fig.3-A shows perfect trajectory tracking while Fig.3-B event tracking, and the latter property is robust. In terms of computation, disturbance rejection may be seen as frequency recognition as in [5]; tracking as synchronization to external events (an insect following a prey).
IV Discussion, Implications, Conclusions
This paper characterized reliability of neurons in terms of average contraction. In excitable networks, contraction was shown to yield regulation to a robustly stable steady state. We propose that regulation can be in turn seen as a generalized dynamical form of computation. As regulation is robust, and spiking is robustly detectable via threshold nonlinearities, this model of computation is robust albeit being analog. Although the results have been developed for the particular case of FHN models, we believe that the main conclusions extend to more general conductance-based models, and that the results of this paper may be the seed for a theory of reliable computation in general neuromorphic systems.
A fundamental property of the proposed model of computation is that regulation is inherently input-selective as so is reliability. Input-selectivity of stability is a sharp departure point from classical control, where properties like synchronization and stability are typically conceived as system properties, independent from, and uniform in, the input, and achieved via high-gain couplings. Here coupling gains define the networks’ behavior, such as the internal models, and play no role in stabilization. Freeing the coupling gains from stabilization comes with several advantages: it avoids the noise amplification bottleneck of high-gain control, opens the door to robust adaptation, and enables the use of weak, synaptic gains that are more suitable to event-based control [26].
The necessity of input-selectivity in neuronal computation is also advocated in neuroscience – as R. Brette observes in [5], “for synchrony to be computationally useful, it must be stimulus dependent”. Moreover, input-selectivity suggests a way to automatically filter relevant input information without higher-level supervision, a paradox of current digital intelligence. In particular, one may think of a redundant parallel interconnection of many networks subject to the same input and whose output are projected downstream via averaging. If the input leads to unreliability, the outputs of the networks are incoherent as they depend on the initial conditions, and they will average to zero: the projected output has maximum entropy and zero mean. If the input induces contraction, all outputs synchronize, and so does the projected output. As a consequence, reliable responses are transmitted downstream due to synchronization whereas unreliable ones are not.
Finally, another consequence of input-selectivity is that it makes regulation inherently multiresolution: single spikes are not contractive events per se, as they visit both contraction and expansion regions. Instead, reliability emerges in the time scale of several spikes, depending on the amount of time spent by the state in the contraction regions. By the same principle, in a longer time interval contraction may be lost, and then regained again in a longer one. This reflects the natural spatio-temporal hierarchy of nervous systems: several spikes form a burst, several bursts activate a muscle, etc.
A further key takeaway of the developed analysis is that contraction, albeit being a form of stability, does not refer to an equilibrium. Indeed, stability of the rest point of the FHN model is not necessary: as demonstrated in [7], the same behavior is observed if the parameters in (1) are chosen so that its equilibrium is unstable. On the contrary, input-dependent reliability would not be possible without instability, which, in turn, is at the core of excitability. As a consequence, the proposed analysis suggests a new non-equilibrium approach to the study of excitable dynamics in terms of contraction that departs from canonical approaches to computational neuroscience based on constant inputs, equilibria, limit cycles, and nullclines [16].
Figure 4: (left) simulation of (11) from 55 random initial conditions; (right) simulation with $u$ changed to $u(t)=\sin(2\omega t)$.
In addition, this interplay between input-selectivity, instability, and contraction enables us to establish an interesting connection to classical linear regulation and robust harmonic rejection in nonlinear systems. Every asymptotically stable LTI system is contractive; yet, contraction is uniform in the input, so as asymptotically stable LTI systems are not a relevant case since they lack input selectivity. Input selectivity is instead observed in forced, undamped linear oscillators, which become (Lagrange [1]) unstable when subject to resonant sinusoidal inputs matching their natural frequency. For the sake of illustration, consider the following equation (the drawn conclusions extend easily to general oscillators)
$\ddot{y}(t)+\omega^{2}y(t)=u(t),\qquad u(t)=\sin(\omega t),$ (11)
describing a linear oscillator with natural frequency $\omega$ forced by a sinusoidal input with the same frequency. Solutions are of the form $y(t)=\alpha\sin(\omega t+\phi)+\beta t\sin(\omega t+\gamma)$, for suitable values of $\alpha,\beta,\phi,\gamma\in\mathbb{R}$ depending on the initial conditions. Clearly, (11) is not contractive being not asymptotically stable. Thus, trajectories $y$ obtained for different initial conditions $y(0)$ and $\dot{y}(0)$ will not converge to one another. Yet, they will synchronize their peaks. Indeed, peaks occur at all times $t^{\star}$ in which $\dot{y}(t^{\star})=0$. Substituting the previous expression and solving for $\cos(\omega t^{\star}+\gamma)$ yields $\cos(\omega t^{\star}+\gamma)=-(\beta\omega t^{\star})^{-1}(\alpha\omega\cos(\omega t^{\star}+\phi)+\beta\sin(\omega t^{\star}+\gamma))$. As the peak times $t^{\star}$ increase to infinity, we thus obtain $\cos(\omega t^{\star}+\gamma)\to 0$. Therefore, independently from the particular initialization, peak times synchronize to the zeros of $\cos(\omega\,\cdot+\gamma)$. Fig. 4 depicts a simulation showing that inputs with frequency $\omega$ induce peaks synchronization, while inputs with a different frequency do not. The role of instability in input-selective synchronization thus shines even more in the LTI case.
While in this LTI case we cannot talk about contraction in the strict sense, the observed behavior has all the elements of the input-selective contraction property of excitable systems studied in this paper. And, indeed, linear oscillators are the very heart of Francis and Wonham’s original internal model principle [10], and their proper embedding in the control loop is both necessary and sufficient for regulation in the LTI case. In addition, as proved in [2, Prop. 3], in the presence of contraction (incremental input-to-state stability), internal models made of LTI oscillators also guarantee robust rejection of harmonics at the embedded frequencies in general nonlinear systems. As sinusoids can be seen as a limit-case of events maximally concentrated in frequency and maximally spread in time – i.e., the presence of an harmonic – this ultimately creates a link to event regulation, which, in these terms, appears as a robust property also in classical nonlinear regulation.